OpenStudy (anonymous):

1. If y = 8, then 5y = ? A. 13 B. 3 C. 40 D. 1.6 2. If x = 9, then 11x = ? A. 20 B. 119 C. 99 D. 911 3. Evaluate the expression for m =2 and n =-3. 3m - 8n A. -25 B. -18 C. 30 D. 7 4. If a = 7, then 9a = ? A. 97 B. 63 C. 79 D. 36 5.Which of the following expressions represents a number m,

OpenStudy (anonymous):

my answer for #1 is a and 2 is a and 3 is c and 4 idk and 5 is b

ganeshie8 (ganeshie8):

#1 is wrong

ganeshie8 (ganeshie8):

#2 is also wrong

OpenStudy (anonymous):

1 is 3

ganeshie8 (ganeshie8):

#1 is 40 = \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

wat about #2 ?

OpenStudy (anonymous):

how u got 40

ganeshie8 (ganeshie8):

u oly said : 1 is 3 ?

ganeshie8 (ganeshie8):

didn't u mean for #1, option C is wat u got ?

ganeshie8 (ganeshie8):

anyways, here is how we evaluate these :- y = 8 5y = 5*y = 5*8 = 40

ganeshie8 (ganeshie8):

you need to multiply... not add.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok wait

ganeshie8 (ganeshie8):

okay

OpenStudy (anonymous):

#1 8x5=40 #2 9x11=99

ganeshie8 (ganeshie8):

right

ganeshie8 (ganeshie8):

wat about #3 ?

OpenStudy (anonymous):

b is 3x(2)-8x(-3) so its 30

ganeshie8 (ganeshie8):

right

OpenStudy (anonymous):

4 is 63

OpenStudy (anonymous):

5 is 13+m

ganeshie8 (ganeshie8):

right right.. looks u mastered these :) next questions im not gona check... il simply say right... :P

ganeshie8 (ganeshie8):

hey for #5, i dont see options ?

OpenStudy (anonymous):

Which of the following expressions represents a number m, multiplied by 13? A. 13m B. 13 + m C. 13 - m D. m – 13

ganeshie8 (ganeshie8):

#5 it should be : 13m

ganeshie8 (ganeshie8):

13+m represents " a number m, ADDED to 13"

OpenStudy (anonymous):

damn your smart what grade r u in

ganeshie8 (ganeshie8):

ahh no, you're smart ! im working... done wid studies lol..